Operations Report – January 8th

Operations Report

Crew 219 Operations Report 08-JAN-2020

Sol: 3

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 17:00

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 59%

Diesel Reading - 73%

Station Propane Reading- 51%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147473,1

Water (static tank) (gal) – 506

Static to loft Pump used – Yes

Water in Green Hab (gal): 200.02

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Diemos rover used: Still in the Work shop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

EVA #6

Starting Hours: 123.5

Beginning Charge: 100%

Ending Hours: 123.7

Ending Charge: 68%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

EVA #6

Starting Hours: 129.3

Beginning Charge: 100%

Ending Hours: 129.5

Ending Charge: 70%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used by Outpost team to take Crew car to
the Work shop, refilled with gas.

CrewCar used and why, where? Crew car was brought to the workshop, is
off site, and was refilled.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: A two- piece helmet had vents that were
stuck aiming downward providing no defogging during EVA #5. The vents
freely rotate within this helmet and cannot be secured. Transformer
for Charging unit 4 is giving very high amperage and voltage exceeding
the limits for the radio batteries. This transformer has been marked
not to be used. Also, during EVA there was a window where EVA radios
were giving a persistent static. The radios are being tested and
troubleshot for any deficiencies in connectivity, battery performance,
and individual radio issues.

Summary of Hab: Pipe froze again today. The heating coil is still
functioning and the pump is still operating properly. The freezing
seems to be in between the exterior wall and the interior wall leading
to the washroom. This was solved using a combined method of warm water
through the feedline, a hair dryer heating the interior and exterior
pipe through the wall.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: Requesting for
Teflon tape for the loft tank feedline and new insulation for the
Static tank to Loft tank pipe.
Copyright © The Mars Society. All rights reserved. | Main Site